# Age-of-Information-Aware Distributed Task Offloading and Resource Allocation in Mobile Edge Computing Networks

### v1

@@ 7 Conclusion
Age-of-Information-Aware Distributed
Task Offloading and Resource Allocation
in Mobile Edge Computing Networks
@@ Appendix B Proof of Theorem 1
    $\sum_{j\in\mathcal{E}}\left(\sum_{i\in\mathcal{M}}\sqrt{D_{\text{ES}(s),ij}}x_{ij}\right)^{2}+\sum_{j\in\mathcal{E}}\left(\sum_{i\in\mathcal{M}}\sqrt{D_{\text{cm},ij}}x_{ij}\right)^{2}$
@@ B.1 Conservation of the Weak Duality Condition
$=\min_{\mathbf{a},\mathbf{b},\mathbf{X}\in\mathcal{F}}\sum_{j\in\mathcal{E}}\left(\sum_{i\in\mathcal{M}}\sqrt{D_{\text{ES}(s),ij}}x_{ij}\right)^{2}$
$~{}~{}~{}~{}+\sum_{j\in\mathcal{E}}\left(\sum_{i\in\mathcal{M}}\sqrt{D_{\text{cm},ij}}x_{ij}\right)^{2}$
@@ B.2 Duality Gap
$\hat{f}=$ $\sum_{j\in\mathcal{E}}\left(\sum_{i\in\mathcal{M}}\sqrt{D_{\text{ES}(s),ij}}x_{ij}\right)^{2}$
$+\sum_{j\in\mathcal{E}}\left(\sum_{i\in\mathcal{M}}\sqrt{D_{\text{cm},ij}}x_{ij}\right)^{2}+\sum_{i\in\mathcal{M}}\sum_{j\in\mathcal{E}}c_{ij}x_{ij}$ (S19)
@@
where $\Delta_{j}^{(1)}=\sqrt{a_{j}}-\left(\sum_{i\in\mathcal{M}}\sqrt{D_{\text{cm},ij}}x_{kj}\right)$ and $\Delta_{j}^{(2)}=\sqrt{b_{j}}-\left(\sum_{i\in\mathcal{M}}\sqrt{D_{\text{ES}(s),ij}}x_{ij}\right)$. Then, we have the optimality gap as below:

### v2

@@ 7 Conclusion
Distributed Task Offloading and Resource Allocation for Latency Minimization in Mobile Edge Computing Networks
@@ Appendix B Proof of Theorem 1
    $\sum_{j\in\mathcal{E}}\left(\sum_{i\in\mathcal{M}}\sqrt{\hat{D}_{\text{ES}(s),ij}}x_{ij}\right)^{2}+\sum_{j\in\mathcal{E}}\left(\sum_{i\in\mathcal{M}}\sqrt{\hat{D}_{\text{cm},ij}}x_{ij}\right)^{2}$
@@ B.1 Conservation of the Weak Duality Condition
$=\min_{\mathbf{a},\mathbf{b},\mathbf{X}\in\mathcal{F}}\sum_{j\in\mathcal{E}}\left(\sum_{i\in\mathcal{M}}\sqrt{\hat{D}_{\text{ES}(s),ij}}x_{ij}\right)^{2}$
$~{}~{}~{}~{}+\sum_{j\in\mathcal{E}}\left(\sum_{i\in\mathcal{M}}\sqrt{\hat{D}_{\text{cm},ij}}x_{ij}\right)^{2}$
@@ B.2 Duality Gap
$\hat{f}=$ $\sum_{j\in\mathcal{E}}\left(\sum_{i\in\mathcal{M}}\sqrt{\hat{D}_{\text{ES}(s),ij}}x_{ij}\right)^{2}$
$+\sum_{j\in\mathcal{E}}\left(\sum_{i\in\mathcal{M}}\sqrt{\hat{D}_{\text{cm},ij}}x_{ij}\right)^{2}+\sum_{i\in\mathcal{M}}\sum_{j\in\mathcal{E}}c_{ij}x_{ij}$ (S19)
@@
where $\Delta_{j}^{(1)}=\sqrt{a_{j}}-\left(\sum_{i\in\mathcal{M}}\sqrt{\hat{D}_{\text{cm},ij}}x_{kj}\right)$ and $\Delta_{j}^{(2)}=\sqrt{b_{j}}-\left(\sum_{i\in\mathcal{M}}\sqrt{\hat{D}_{\text{ES}(s),ij}}x_{ij}\right)$. Then, we have the optimality gap as below: